Two tugs are pulling a vessel in water. Tug A is at an angle of 25º to the direction of travel and tug B is at an angle of 10º to the direction of travel. A resultant force of 150kN is produced by the two tugs. Calculate the force in the tugs ropes. 


1
Expert's answer
2021-01-21T09:38:42-0500

"f_1\cdot\cos10°+f_2\cdot\cos25°=150"


"f_1\cdot\sin10°-f_2\cdot\sin25°=0"


solving this system of linear equations, we obtain


"f_1\approx110.5(N)" and "f_2\approx45.4(N)" . Answer